OpenStudy (anonymous):

(g+9)/(g-3) ÷ (3g+27)/(g-5). The answer I got is 1/g(3g-24)+45. Is this correct?

OpenStudy (anonymous):

I got (g-5)/3(g-3)

OpenStudy (turingtest):

it should be that times (g+9)/(g-3) which is(g-5)(g+9)/[3(g-3)^2] you could foil it out but I don't see much to simplify...

OpenStudy (anonymous):

LOL, boy this is confusing....

OpenStudy (anonymous):

start with \[\frac{g+9}{g-3}\times \frac{g-5}{3g+27}\] and then look to factor as much as you can

OpenStudy (turingtest):

exactly

OpenStudy (anonymous):

not much to factor really, just get \[\frac{g+9}{g-3}\times \frac{g-5}{3(g+9)}=\frac{\cancel{g+9}}{g-3}\times \frac{g-5}{3\cancel{g+9}}=\frac{g-5}{3(g-3)}\]

OpenStudy (anonymous):

So what exactly does that translate to because I am confused.

OpenStudy (anonymous):

(g-5)/(g-3) or (g-5)(g+9)/3(g-3)^2...either one looks like it could be right.

OpenStudy (turingtest):

satellite was right and did it the simplest way. I think I made a mistake.

OpenStudy (turingtest):

oh yeah, I should have gotten (g-5)(g+9)/[3(g+9)(g-3)] which simplifies to (g-5)/[3(g-3)]